OpenStudy (anonymous):

A greenhouse that specializes in growing bell peppers is divided into sections. The number of plants in each section depends on the number of sprinklers in that section. In a section with x sprinklers, there are 3x(x + 3) red bell pepper plants and (x + 4)2 yellow bell pepper plants. Which simplified expression represents the total number of red and yellow bell pepper plants in a section with x sprinklers? answer choices: 3x2 + 17x + 16 4x2 + 17x + 16 3x4 + 17x2 + 16 4x4 + 17x2 + 16

ganeshie8 (ganeshie8):

red bell pepper plants : 3x(x+3) = 3x^2+9x yellow pepper plants : (x+4)^2 = x^2 + 8x + 16 read + yellow pepper plants = ?

ganeshie8 (ganeshie8):

you just need to add red and yellow. thats it right ?

OpenStudy (anonymous):

So would it be 4x2 + 17x + 16?

ganeshie8 (ganeshie8):

thats right !

OpenStudy (anonymous):

Thank you! @ganeshie8

ganeshie8 (ganeshie8):

np. .